\newpage
\begin{thebibliography}{AGKS}

\bibitem{AKNS}
\textsc{M.J. Ablowitz, D.J. Kaup, A.C. Newell, H. Segur},  \textit{The
inverse
scattering transform - Fourier analysis for nonlinear
problems,}
Stud. Appl. Math. \textbf{53} (1974) 249--315.

\bibitem{AS}
\textsc{M.J. Ablowitz, H. Segur},  \textit{ Solitons and the inverse
scattering
transform,}  SIAM Stud. Appl. Math. \textbf{4},
Philadelphia,
1981.



\bibitem{AGKLS1}
\textsc{D. Alpay, I. Gohberg, M.A. Kaashoek, L. Lerer,  A. Sakhnovich,} \textit{ 
Krein systems},  in: Oper. Theory Adv. Appl. {\bf 191} (2009), Birkh\"auser, 19--36.

\bibitem{AGKLS2}
\textsc{D. Alpay, I. Gohberg, M.A. Kaashoek, L. Lerer, A. Sakhnovich,}
{\it Krein systems and canonical systems on a finite interval: accelerants
with a jump discontinuity at the origin and  continuous potentials},
Integral Equations Operator Theory 
{\bf 68}:1 (2010) 115--150.

\bibitem{AD0}
\textsc{D.Z. Arov,  H. Dym,} 
 \textit{Direct and inverse problems for differential systems connected with Dirac systems and related factorization problems,}
 Indiana Univ. Math. J.  \textbf{54}:6  (2005) 1769--1815. 

\bibitem{AD}
\textsc{D.Z. Arov, H. Dym, } \textit{
$J$-contractive matrix valued functions and related topics.} 
Encyclopedia of Mathematics and its Applications \textbf{116},
Cambridge University Press, Cambridge, 2008.

%\bibitem{AFK0} D.Z. Arov, B. Fritzsche, B. Kirstein, Completion problems for $j\sb {pq}$-inner functions I, Integral Equations Operator Theory 16 (1993) 155--185.

%\bibitem{AFK} D.Z. Arov, B. Fritzsche, B. Kirstein, Completion problems for $j\sb {pq}$-inner functions II,  Integral Equations Operator Theory 16 (1993) 453--495.



%\bibitem{B2} J.A. Ball, Linear systems, operator model theory and scattering: multivariable generalizations, Fields Inst. Commun. 25 (2000) 151--178.


%\bibitem{BC} R. Beals, R.R.  Coifman, Scattering and inverse scattering for first order systems,  Comm. Pure Appl. Math. 37 (1984) 39--90.

%\bibitem{BDT} R. Beals, P. Deift, C. Tomei,  Direct and inverse scattering on the line, in: Mathematical Surveys and Monographs, Vol. 28, AMS, Providence,  1984.


\bibitem{CS} 
\textsc{K. Chadan, P.C.  Sabatier,}  \textit{Inverse problems in quantum scattering theory,} Springer, New York, 1989.

\bibitem{Ci}
\textsc{J.L. Cie\'sli\'nski,}  \textit{Algebraic construction of the Darboux matrix revisited,} J. Phys. A \textbf{42}:40 (2009),  404003, 40 pp.


\bibitem{CG1} 
\textsc{S. Clark, F.  Gesztesy,} \textit{
Weyl-Titchmarsh $M$-function asymptotics, local uniqueness results, trace formulas, and Borg-type theorems for Dirac
operators,} Trans. Amer. Math. Soc. \textbf{354} (2002)
3475--3534.




\bibitem{CG2}  
\textsc{S. Clark, F.  Gesztesy,} \textit{ 
On Self-adjoint and J-self-adjoint Dirac-type Operators: A Case Study},  Contemp. Math. {\bf 412} (2006)  103--140.


%\bibitem{CGZ} S. Clark,  F.  Gesztesy  and M. Zinchenko, {\it Weyl-Titchmarsh theory and Borg-Marchenko-type  uniqueness results for CMV operators with matrix-valued Verblunsky coefficients},  Oper. Matrices  {\bf 1}:4  (2007),  535--592.


%\bibitem{CGZ2} S. Clark,  F.  Gesztesy  and M. Zinchenko, {\it  Borg-Marchenko-type uniqueness results for CMV operators}, Skr. K. Nor. Vidensk. Selsk.  2008,  {\bf 1} (2008), 1--18.


\bibitem{cr}
\textsc{M. M. Crum,} \textit{ Associated Sturm--Liouville systems}, Quart. J.
Math. Oxford (2) {\bf 6} (1955) 121--127.

\bibitem{de}
\textsc{P.A. Deift,} \textit{ Applications of a commutation formula}, Duke
Math. J. {\bf 45} (1978) 267--310.



%\bibitem{DFK} V.K. Dubovoj, B. Fritzsche, B. Kirstein, Matricial version of the classical Schur problem, in: Teubner-Texte zur Mathematik [Teubner Texts in Mathematics], Vol. 129, B. G. Teubner Verlagsgesellschaft mbH, Stuttgart, 1992.


\bibitem{FT}
\textsc{L.D. Faddeev, L.A. Takhtajan,}   \textit{Hamiltonian methods in the theory of solitons,} Springer, New York, 1986.

%\bibitem{FK1} B. Fritzsche, B. Kirstein, M. Mosch, On block completion problems for Arov-normalized $j\sb {qq}$-$J\sb q$-elementary factors, Lin. Alg. Appl. 346 (2002) 273--291.

\bibitem{FKRS}
\textsc{B. Fritzsche, B. Kirstein, I.Ya. Roitberg, A.L. Sakhnovich,} \textit{ Weyl matrix functions and 
inverse problems for discrete Dirac type self-adjoint system: explicit and general 
solutions}, Operators and Matrices \textbf{2} (2008)  201--231.

\bibitem{FKRS3}
\textsc{B. Fritzsche, B. Kirstein, I.Ya. Roitberg, A.L. Sakhnovich, }
\textit{ Weyl theory  and inverse problem for a general-type 
Dirac system  with rectangular matrix potential}, Preprint.

\bibitem{FKS0}
\textsc{B. Fritzsche, B. Kirstein, and A.L. Sakhnovich,} \textit{ Completion
problems and scattering problems for Dirac type differential
equations with singularities},  J. Math. Anal. Appl. \textbf{317}
(2006) 510--525.

\bibitem{FKS1}
\textsc{B. Fritzsche, B. Kirstein,   A.L. Sakhnovich,} \textit{ 
Semiseparable integral operators and explicit solution of an 
inverse problem for the skew-self-adjoint Dirac-type system,} 
Integral Equations Operator Theory {\bf 66} (2010) 231--251.

\bibitem{FKS2}
\textsc{B. Fritzsche, B. Kirstein,   A.L. Sakhnovich,} \textit{
Weyl Functions of Generalized Dirac Systems: Integral Representation,
the Inverse Problem and Discrete Interpolation},
J. Anal. Math. to appear.

\bibitem{fg}
\textsc{F. Gesztesy,} {\textit A complete spectral characterization of the
double commutation method}, J. Funct. Anal. {\bf 117} (1993) 401--446.


\bibitem{GeH}
\textsc{F.~Gesztesy, H.~Holden,}  \textit{ Soliton equations and their
algebro-geometric solutions,} Cambridge Studies in Advanced
Mathematics \textbf{79}, Cambridge University Press, Cambridge, 2003.





\bibitem{GSS}
\textsc{F. Gesztesy, W.  Schweiger, B. Simon,} \textit{ Commutation
methods applied
to the mKdV-equation,} Trans. Amer. Math. Soc. \textbf{324} 
(1991) 465--525.

%\bibitem{GeSi0} F. Gesztesy  and B. Simon,   {\it On local Borg-Marchenko uniqueness results},  Commun. Math. Phys. {\bf 211} (2000), 273--287.


%\bibitem{GeSi} F. Gesztesy and B.  Simon, {\it  A new approach to inverse spectral theory, II,  General real potentials and the connection to the spectral measure},  Ann. of Math. (2)  {\bf 152}:2  (2000),  593--643.

\bibitem{gt}
\textsc{F. Gesztesy, G. Teschl,} \textit{ On the double  commutation method},  Proc. Amer. Math. Soc. {\bf 124} (1996) 1831--1840.

\bibitem{GKS5}
\textsc{I. Gohberg, M.A. Kaashoek, A.L. Sakhnovich,} \textit{ Bound states
for canonical systems on the half and full line: explicit
formulas}, Integral Equations  Operator Theory {\bf 40}:3
(2001) 268--277.


\bibitem{GKS6}
\textsc{I. Gohberg, M.A. Kaashoek,  A.L. Sakhnovich,} \textit{ Scattering
problems for a canonical system with a pseudo-exponential
potential},  Asymptotic Analysis {\bf 29}:1 (2002) 1--38.


%\bibitem{gohkol} I.~Gohberg and I.~Koltracht, Numerical solution of integral equations, fast algorithms and {K}rein--{S}obolev equations, \emph{Numer. Math.} \textbf{47} (1985), 237--288.

%\bibitem{GoKr} I.Gohberg  and  M.G.Krein,  Transl.  of  math.  monographs  {\bf 24}, Providence,  Rhode Island,  1970.

\bibitem{Gu}
\textsc{C.H.~Gu, H.~Hu,  Z.~Zhou,} \textit{Darboux transformations in
integrable systems,} Springer Verlag, 2005.

\bibitem{KFA}
\textsc{R.E. Kalman, P. Falb, M. Arbib,} \textit{ Topics in mathematical system
theory}, McGraw-Hill, New York, 1969.



\bibitem{KoSaTe}
 \textsc{A.~Kostenko, A.~Sakhnovich,  G.~Teschl,} \textit{
Commutation Methods for Schr\"odinger Operators with Strongly Singular Potentials}, arXiv:1010.4902 (Mathematische Nachrichten to appear).



\bibitem{Kre1}
\textsc{M.G. Krein,} \textit{ Continuous analogues of propositions on polynomials
orthogonal on the unit circle} (Russian), Dokl. Akad. Nauk SSSR
{\bf 105} (1955) 637--640.

\bibitem{kr57}
\textsc{M.G. Krein,} \textit{ On a continual analogue of a Christoffel formula from the theory of orthogonal polynomials} (Russian),
Dokl. Akad. Nauk SSSR (N.S.)  {\bf 113} (1957) 970--973. 


\bibitem{K}
\textsc{M.G. Krein,} \textit{Topics in differential and integral
equations and
operator theory,}  Oper. Theory Adv. Appl. \textbf{7},
Birkh\"auser, Basel-Boston, 1983.

\bibitem{LR}
\textsc{P. Lancaster, L. Rodman,}
\textit{ Algebraic Riccati equations}, Clarendon Press, Oxford,
1995.



%\bibitem{LaWo} M. Langer and H. Woracek, {\it   A local inverse spectral theorem for Hamiltonian systems},  ASC Report No. 43/2009, http://www.asc.tuwien.ac.at/preprint/2009/asc43x2009.pdf.


\bibitem{LS}
\textsc{B.M. Levitan, I.S.  Sargsjan,} \textit{ Sturm-Liouville and
Dirac
operators,}  Mathematics and its Applications \textbf{59}, Kluwer,
Dordrecht, 1990.


%\bibitem{L} M.S. Liv\v{s}ic,  On a class of linear operators in Hilbert space, in:   Amer. Math. Soc. Transl. (2), Vol. 13, 85-103, 1960.

\bibitem{Mar}
\textsc{V.A. Marchenko,}  \textit{ Sturm-Liouville operators and applications,}
Oper. Theory Adv. Appl. {\bf 22}, Birkh\"auser, Basel, 1986.

\bibitem{Mar2}
\textsc{V.A.~Marchenko,} \textit{Nonlinear equations and operator algebras,}
Reidel Publishing Co., Dordrecht, 1988.

\bibitem{MS}
\textsc{V.B.~Matveev, M.A.~Salle,} \textit{ Darboux transformations and
solitons,} Springer Verlag, Berlin, 1991.



%\bibitem{WP}  R.E.A.C. Paley and N. Wiener, {\it  Fourier transforms in the complex domain},   American Mathematical Society Colloquium Publications {\bf 19}, Amer. Math. Soc., Providence, RI, 1987.


\bibitem{P}  
\textsc{V.P. Potapov,} \textit{The multiplicative structure of $J$-contractive
matrix functions,} Amer. Math. Soc. Transl. {\bf 15} (1960) 131--243.



%\bibitem{SaA0} A.L. Sakhnovich, {\it Asymptotics of spectral functions of an $S$-colligation},  Soviet Math. (Iz. VUZ) {\bf 32} (1988), 92-105.


\bibitem {SaA1}
\textsc{A.L. Sakhnovich,}  \textit{ A nonlinear Schr\"odinger equation on the
semiaxis and a related inverse  problem},  Ukrain. Math. J. \textbf{
42}:3 (1990) 316--323.

\bibitem {SaA2}
\textsc{A.L. Sakhnovich,} \textit{ The Goursat problem for the sine-Gordon
equation and the inverse spectral problem},
 Russ. Math. Iz. VUZ \textbf { 36}:11 (1992) 42--52.
 
\bibitem{ALS94}
\textsc{ A.L.~Sakhnovich,} \textit{ Dressing procedure for solutions of nonlinear
equations and the method of operator identities.} { Inverse
Problems},  { 10} (1994) 699--710.

\bibitem{SaA21}
\textsc{A.L. Sakhnovich,} \textit{Inverse spectral problem related to the
$N$-wave equation},   in:  {  M.G. Krein volume}, Oper. Theory Adv. Appl. \textbf{  117} 
(2000),  Birkh\"auser, Basel,  323--338.



\bibitem{SaA3}
\textsc{A.L. Sakhnovich,} \textit{Dirac type and canonical  systems:
spectral and
Weyl-Titchmarsh fuctions, direct and inverse problems,}
Inverse
Problems  \textbf{18} (2002) 331--348.



\bibitem{SaA7}
\textsc{A.L. Sakhnovich,} \textit{ Second harmonic generation: Goursat problem
on the semi-strip, Weyl functions and explicit solutions}, Inverse
Problems \textbf{ 21}:2 (2005) 703--716.

\bibitem{ALS10}
\textsc{A.L. Sakhnovich,} {\textit On the GBDT version of the B\"acklund-Darboux transformation and its applications to the
linear and nonlinear equations and spectral theory},
Mathematical Modelling of Natural Phenomena {\bf 5}:4 (2010) 340--389.



%\bibitem{ASAK} A.L. Sakhnovich, A.A. Karelin, J. Seck-Tuoh-Mora, G. Perez-Lechuga and M.  Gonzalez-Hernandez, {\it  On explicit inversion of a subclass of operators with $D$-difference kernels  and Weyl theory of the corresponding canonical systems}, Positivity {\bf 14}  (2010), 547-564. 


%\bibitem{SaL0} L.A. Sakhnovich, {\it Spectral analysis of Volterra's operators defined in the space of vector-functions $L_m^2(0,l)$}, Ukr. Mat. Zh.  {\bf 16}:2 (1964),  259-268.

\bibitem{SaL1}
\textsc{L.A. Sakhnovich,} \textit{On  the  factorization  of  the 
transfer matrix
function,}  Sov. Math. Dokl. \textbf{17} (1976) 203--207.


\bibitem{SaL2}
\textsc{L.A. Sakhnovich,} \textit{Factorisation  problems  and 
operator
identities,}   Russian
 Math. Surv. \textbf{41} (1986) 1--64.

\bibitem{SaL2'}
\textsc{L.A. Sakhnovich,} \textit{ Integral equations with difference kernels
on finite intervals}, Oper. Theory Adv. Appl. {\bf 84},
Birkh\"auser, Basel-Boston-Berlin, 1996.

\bibitem{SaL3}
\textsc{L.A. Sakhnovich,} \textit{Spectral theory of canonical
differential
systems. Method of operator identities,}  Oper.
Theory Adv.
Appl. \textbf{107}, Birkh\"auser, Basel-Boston, 1999.

\bibitem{SaL4}
\textsc{L.A. Sakhnovich, } \textit{
Matrix finite-zone Dirac-type equations,} 
 J. Funct. Anal.  \textbf{193}:2  (2002),  385--408.
 
%\bibitem{Si}  B. Simon,  {\it    A new approach to inverse spectral theory, I, Fundamental formalism},  {   Ann. of Math.} { \bf 150} (1999), 1029--1057. 

\bibitem{T1}
\textsc{G. Teschl,} \textit{
Jacobi operators and completely integrable nonlinear lattices,}
Mathematical Surveys and Monographs,  Amer. Math. Soc., Rhode Island, 2000.

\bibitem{T2}
\textsc{G. Teschl,} \textit{ Mathematical Methods in Quantum Mechanics; With Applications to Schr\"odinger Operators},
Graduate Studies in Mathematics, Amer. Math. Soc., Rhode Island, 2009.

%\bibitem{V} V. Vinnikov, Commuting operators and function theory on a Riemann surface, in: Math. Sci. Res. Inst. Publ., Vol. 33, 445--476, Cambridge Univ. Press, Cambridge, 1998.

\bibitem{ZM}
\textsc{V.E.~Zakharov, A.V.~Mikhailov,} \textit{ On the integrability of
classical spinor models in two-dimensional space-time.} Comm.
Math. Phys. \textbf{74} (1980) 21--40.


\bibitem{ZS}
\textsc{V.E. Zaharov, A.B. Shabat,} \textit{On soliton interaction in
stable media,}
JETP \textbf{64} (1973) 1627--1639.

%\bibitem{Zh} X. Zhou, Inverse scattering transform for systems with rational spectral dependence, J. Diff. Eqs.  115 (1995)  277--303.

 
\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{cite}
\newsymbol \blackbox 1004
\newcommand{\eh}{\hfill}\newlength{\sperr}

\newenvironment{proof}{{\settowidth{\sperr}{\bf\rm
Proof}%
\par\addvspace{0.3cm}\noindent\parbox[t]{1.3\sperr}
{\bf\rm P\eh r\eh o\eh o\eh f\eh }%
}}{\nopagebreak\mbox{}
$\blackbox$\par\addvspace{0.3cm}}

\def\nn{\nonumber}
\def\a{\alpha}
\def\b{\beta}

\def\Lam{\Lambda}
\def\s{\sigma}
\def\la{\lambda}
\def\om{\omega}

\def\S{\Sigma}
\def\t{\theta}
\def\Up{\Upsilon}
\def\vp{\varphi}
\def\vt{\vartheta}
\def\ve{\varepsilon}
\def\wh{\widehat}

\def\ov{\overline}

\def\BC{{\mathbb C}}
\def\BR{{\mathbb R}}
\def\BN{{\mathbb N}}
\def\clp{{\mathcal P}}
\def\cla{{\mathcal A}}

\def\clh{{\mathcal H}}

\def\cln{{\mathcal N}}

\def\cld{{\mathcal D}}

\def\im{{\rm Im\ }}

\newcommand{\E}{\mathrm{e}}
\newcommand{\I}{\mathrm{i}}
\def\mf{\mathfrak}

\newtheorem{Pa}{Paper}[section]
\newtheorem{Tm}[Pa]{{\bf Theorem}}

\newtheorem{Cy}[Pa]{{\bf Corollary}}
\newtheorem{Rk}[Pa]{{\bf Remark}}

\newtheorem{Dn}[Pa]{{\bf Definition}}
\newtheorem{Pn}[Pa]{{\bf Proposition}}

\title{Weyl theory and explicit solutions of direct and inverse
problems for a Dirac system with rectangular matrix potential  }

\author{B. Fritzsche, B. Kirstein, I.Ya. Roitberg, A.L. Sakhnovich}

\date{}
\begin{document}
\maketitle

\begin{abstract} A non-classical Weyl theory is developed for
Dirac systems with rectangular matrix potentials. The notion
of the Weyl function is introduced and the corresponding
direct problem is treated. Furthermore, explicit solutions of the 
direct and inverse problems are obtained for the case of rational
Weyl matrix functions.

\end{abstract}

{MSC(2010):}  4B20, 34L40, 15A15, 93B15.

Keywords:  {\it Weyl function, Weyl theory, Dirac system, rectangular matrix potential, direct problem, inverse problem, 
pseudo-exponential potential, explicit solution, rational matrix function, contractive matrix function, realization.} 

\section{Introduction} \label{intro}
\setcounter{equation}{0}
Consider  
self-adjoint Dirac-type (also called Dirac, ZS or AKNS)
system, which is
a classical matrix differential equation:
\begin{equation}       \label{1.1}
\frac{d}{dx}y(x, z )=\I (z j+jV(x))y(x,
z ) \quad
(x \geq 0),
\end{equation}
\begin{equation}   \label{1.2}
j = \left[
\begin{array}{cc}
I_{m_1} & 0 \\ 0 & -I_{m_2}
\end{array}
\right], \hspace{1em} V= \left[\begin{array}{cc}
0&v\\v^{*}&0\end{array}\right],
 \end{equation}
where  $I_{m_k}$ is the $m_k \times
m_k$ identity
matrix and $v(x)$ is an $m_1 \times m_2$ matrix function,
which is called the potential of system. 
Dirac-type systems   are very well-known in
mathematics and
applications (see, for instance, books
\cite{AD, CS, K, LS, SaL3},
recent publications \cite{AGKLS1, AGKLS2, AD0,  CG1, CG2, FKS1, FKS2, SaL4},
and numerous references therein).
 The
name ZS-AKNS
is caused by the fact that system (\ref{1.1}) is an
auxiliary
linear system for many important nonlinear integrable
wave
equations and as such it was studied, for instance, in
\cite{AKNS, AS, FT, GSS, SaA7, ZS}.
The Weyl and spectral theory of self-adjoint Dirac systems, where $m_1=m_2$,
was treated, for instance, in \cite{AD0, Kre1, CG1, GKS6, LS, SaA3,  SaL3}
(see also various references therein). The "non-classical" Weyl theory for the
equally important  case $m_1\not= m_2$ 
and related questions are the subject of this paper.  

In Section \ref{DP} we treat the direct problem for the general-type Dirac system,
that is, system \eqref{1.1}, where the potential $v$ is locally summable.
A definition of the non-expansive generalized Weyl function is given,
its existence and uniqueness are proved, and some basic properties are
studied.

In Section \ref{dies} we consider Dirac systems with  so called
{\it generalized pseudo-exponential} potentials (see Definition \ref{Dnpsex}).
Direct and inverse problems for such systems are solved there explicitly.

As usual, $\BN$ stands for the set of natural numbers, $\BR$ stands for the real axis,
$\BC$ stands for the complex plain, and
$\BC_+$ for the open upper
semi-plane. If $a\in \BC$, then $\ov a$ is its complex conjugate.  The notation $\im$
is used for image.
An $m_2 \times m_1$ matrix $\a$ is said to be non-expansive, 
if $\a^*\a \leq I_{m_1}$ (or, equivalently, if $\a\a^* \leq I_{m_2}$).

We put $m_1+m_2=:m$. The fundamental solution of system
\eqref{1.1} is denoted by $u(x,z)$, and this solution is normalized by the condition
\begin{align}&      \label{1.3}
u(0,z)=I_m.
\end{align}
\section{Direct problem} \label{DP}
\setcounter{equation}{0}
We consider Dirac system \eqref{1.1} on the semi-axis $x \in [0,\, \infty)$
and assume that $v$ is measurable and locally summable, that is, summable
on all the finite intervals.
In a way, which is similar, for instance, to the non-classical problem treated in \cite{SaA21}
we shall use M\"obius transformations and matrix balls to solve the direct problem for Dirac
system. 

Introduce a class  of nonsingular $m \times m_1$ matrix functions 
$\clp(z)$ with property-$j$, which are an immediate analog of the classical pairs
of parameter matrix functions. Namely, the matrix functions 
$\clp(z)$ are meromorphic in $\BC_+$ and satisfy
(excluding, possibly, a discrete set of points)
the following relations
\begin{align}\label{2.1}&
\clp(z)^*\clp(z) >0, \quad \clp(z)^* j \clp(z) \geq 0 \quad (z\in \BC_+).
\end{align}
\begin{Dn} \label{set}
The set $\cln(x,z)$ of M\"obius transformations is the set of values at $x, \,z$ 
of matrix functions
\begin{align}\label{2.2}&
\vp(x,z,\clp)=\begin{bmatrix}
0 &I_{m_2}
\end{bmatrix}u(x,z)^{-1}\clp(z)\Big(\begin{bmatrix}
I_{m_1} & 0
\end{bmatrix}u(x,z)^{-1}\clp(z)\Big)^{-1},
\end{align}
where $\clp(z)$ are nonsingular  matrix functions 
 with property-$j$. 
 \end{Dn}
 \begin{Pn} \label{PnW1} Let Dirac system \eqref{1.1} on $[0, \, \infty)$
 be given and assume that $v$ is  locally summable.
 Then the sets $\cln(x,z)$
 are well-defined. There is a unique matrix function
 $\vp(z)$ in $\BC_+$ such that
\begin{align}&      \label{2.3}
\vp(z)=\bigcap_{x<\infty}\cln(x,z).
\end{align} 
This function is analytic and non-expansive.
 \end{Pn}
\begin{proof}.
It is immediate from \eqref{1.1} that
\begin{align}&      \label{2.4}
\frac{d}{dx}\big(u(x,z)^*ju(x,z)\big)=\I (z-\ov z)u(x,z)^*u(x,z)<0, \quad z \in \BC_+.
\end{align}
According to \eqref{1.3} and \eqref{2.4} we have
\begin{align}&      \label{2.5}
\mathfrak{A}(x,z)=\{\mathfrak{A}_{ij}(x,z)\}_{i,j=1}^2:= u(x,z)^*ju(x,z)\leq j,
\quad z \in \BC_+,
\end{align}
where $\mathfrak{A}$ is partitioned into four blocks so that $\mathfrak{A}_{ii}$
is an $m_i \times m_i$ matrix function ($i=1,2$).
Inequality \eqref{2.5} yields
\begin{align}&      \label{2.6}
\big(u(x,z)^*\big)^{-1}ju(x,z)^{-1}\geq j.
\end{align}
Thus, we get 
\begin{align}&      \label{2.7}
\det \Big(\begin{bmatrix}
I_{m_1} & 0
\end{bmatrix}u(x,z)^{-1}\clp(z)\Big)\not= 0,
\end{align}
and so $\cln$ is well-defined via \eqref{2.2}. Indeed, if \eqref{2.7} does not
hold, there is a vector $f \in \BC^{m_1}$ such that
\begin{align}&      \label{2.8}
\begin{bmatrix}
I_{m_1} & 0
\end{bmatrix}j u(x,z)^{-1}\clp(z)f=\begin{bmatrix} I_{m_1} & 0
\end{bmatrix}u(x,z)^{-1}\clp(z)f= 0, \quad f\not=0.
\end{align}
By \eqref{2.1} and \eqref{2.6} the subspace $\im \big(u(x,z)^{-1}\clp(z)\big)$
is a maximal $j$-non-negative subspace. Clearly $\im \big(\begin{bmatrix}
I_{m_1} & 0
\end{bmatrix}^*\big)$ is a maximal $j$-nonnegative subspace too. Therefore
\eqref{2.8} implies $u(x,z)^{-1}\clp(z)f \in \im \big(\begin{bmatrix}
I_{m_1} & 0
\end{bmatrix}^*\big)$. But then it follows from
$$
\begin{bmatrix} I_{m_1} & 0
\end{bmatrix}u(x,z)^{-1}\clp(z)f= 0
$$
that $f=0$, which contradicts the inequality in \eqref{2.8}.

Next, rewrite \eqref{2.2} in the equivalent form
\begin{align}&      \label{2.9}
\begin{bmatrix}
I_{m_1} \\ \vp(x,z,\clp)
\end{bmatrix}=u(x,z)^{-1}\clp(z)\Big(\begin{bmatrix}
I_{m_1} & 0
\end{bmatrix}u(x,z)^{-1}\clp(z)\Big)^{-1}.
\end{align}
In view of \eqref{2.1}, \eqref{2.9}, and of the definition of 
$\mathfrak{A}$ in \eqref{2.5},
formula  
\begin{align}&      \label{2.10}
\wh \vp(z) \in \cln(x,z)
\end{align}
is equivalent to
\begin{align}&      \label{2.11}
\begin{bmatrix}
I_{m_1} & \wh \vp(z)^*
\end{bmatrix}\mathfrak{A}(x,z)\begin{bmatrix}
I_{m_1} \\ \wh \vp(z)
\end{bmatrix} \geq 0.
\end{align}
In a standard way, using formula \eqref{2.4} and the equivalence of \eqref{2.10} and \eqref{2.11}, we get
\begin{align}&      \label{2.16}
\cln(x_1,z) \subset \cln(x_2,z)  \quad {\mathrm{for}} \quad x_1>x_2.
\end{align}
Moreover, \eqref{2.11} at $x=0$ means that
\begin{align}&      \label{2.17}
\cln(0,z) =\{\wh \vp(z): \, \wh \vp(z)^* \wh \vp(z)\leq I_{m_1}\}.
\end{align}
By Montel's theorem, formulas \eqref{2.16} and \eqref{2.17}
imply the existence of an analytic and non-expansive matrix function
$\vp(z)$ such that
\begin{align}&      \label{2.18}
\vp(z)\in \bigcap_{x<\infty}\cln(x,z).
\end{align} 
Indeed, because of \eqref{2.16} and \eqref{2.17} we see that the set 
of functions $\vp(x,z,\clp)$ of the form \eqref{2.2} is uniformly bounded
in $\BC_+$. So, Montel's theorem is applicable and there is an analytic
matrix function, which we denote by $\vp_{\infty}(z)$ and which is a uniform limit
of some sequence
\begin{align}&      \label{M1}
\vp_{\infty}(z)=\lim_{i \to \infty} \vp(x_i,z,\clp_i) \quad (i \in \BN, \quad x_i \uparrow, \quad \lim_{i \to \infty}x_i=\infty)
\end{align} 
on all the bounded and closed subsets of $\BC_+$. Since $x_i \uparrow$
and equalities \eqref{2.9} and \eqref{2.16} hold, it follows that the matrix functions
\begin{align}&      \nn
\clp_{ij}(z):=u(x_i,z)\begin{bmatrix}
I_{m_1} \\ \vp(x_j,z,\clp_j) 
\end{bmatrix} \quad (j \geq i)
\end{align} 
satisfy relations \eqref{2.1}. Therefore, using \eqref{M1} we derive that \eqref{2.1} holds for
\begin{align}&      \nn
\clp_{i, \infty}(z):=u(x_i,z)\begin{bmatrix}
I_{m_1} \\ \vp_{\infty}(z) 
\end{bmatrix},
\end{align} 
which, in its turn, implies that 
\begin{align}&      \label{M2}
\vp_{\infty}(z)\in\cln(x_i,z).
\end{align} 
Since \eqref{M2} holds for all $i\in \BN$,
we see that \eqref{2.18} is true for $\vp(z)=\vp_{\infty}(z)$.

Now, let us show  that $\cln$ is a matrix ball.  It follows from
\eqref{2.5} that $\mf{A}_{22}<-I_{m_2}$. Moreover, it follows 
from \eqref{2.4} and \eqref{2.5} that $\frac{d}{dx}\mf{A} \leq \I(\ov z -z)\mf{A}\leq
 \I(\ov z -z)j$.
Taking into account the inequalities above, we derive
\begin{align}&      \label{2.12}
- \mathfrak{A}_{22}(x,z) \geq \big(1+\I(\ov z -z)x\big)I_{m_2}.
\end{align}
Note also that \eqref{2.5} implies $\mf{A}(x,z)^{-1} \geq j$ 
for $z\in \BC_+$ (see \cite{P}). Thus, we get
\begin{align}&      \label{2.13}
\big(\mathfrak{A}^{-1}\big)_{11}=\big(\mathfrak{A}_{11}-
\mf{A}_{12}\mf{A}_{22}^{-1}\mf{A}_{21}\big)^{-1} \geq I_{m_1}.
\end{align}
As $- \mf{A}_{22}>0$ the square root $\Up=\big(- \mf{A}_{22}\big)^{1/2}$
is well-defined and we rewrite \eqref{2.11} in the form
\begin{align}&      \nn
\mathfrak{A}_{11}-
\mf{A}_{12}\mf{A}_{22}^{-1}\mf{A}_{21}-\big(\wh \vp^*\Up-\mf{A}_{12}\Up^{-1}\big)
\big(\Up \wh \vp-\Up^{-1}\mf{A}_{21}\big) \geq 0, \end{align}
where $\mf{A}_{12}=\mf{A}_{21}^*$. Equivalently, we have
\begin{align}&      \label{2.14}
\wh \vp =\rho_l \om \rho_r-\mf{A}_{22}^{-1}\mf{A}_{21}, \quad \om^*\om \leq I_{m_2},
\\ &      \label{2.15}
 \rho_l:=\Up^{-1}=\big(- \mf{A}_{22}\big)^{-1/2},
\quad \rho_r:=(\mathfrak{A}_{11}-
\mf{A}_{12}\mf{A}_{22}^{-1}\mf{A}_{21}\big)^{1/2}.
\end{align}
Here $\om$ is an $m_2 \times m_1$ matrix function.
Since \eqref{2.10} is equivalent to \eqref{2.14}, the sets
$\cln(x,z)$ (where the values of $x$ and $z$ are fixed) are matrix balls,
indeed.
According to \eqref{2.12}, \eqref{2.13}, and \eqref{2.15}
the next formula holds:
\begin{align}&      \label{2.19}
\rho_l(x,z) \to 0 \quad (x \to \infty), \quad \rho_r(x,z) \leq I_{m_1}.
\end{align}

Finally, relations \eqref{2.18}, \eqref{2.14}, and \eqref{2.19}
imply \eqref{2.3}.

\end{proof}
In view of Proposition \ref{PnW1}
we define the Weyl function similar to the canonical systems 
case \cite{SaL3}. 
\begin{Dn} \label{defWeyl} The Weyl-Titchmarsh (or simply Weyl) function of Dirac system \eqref{1.1} on $[0, \, \infty)$,
where  potential $v$ is locally summable, is the function $\vp$
given by \eqref{2.3}.
\end{Dn}
By Proposition  \ref{PnW1} the Weyl-Titchmarsh function always exists.
Clearly, it is unique.
 
\begin{Cy} \label{CyW2} Let the conditions of Proposition \ref{PnW1} hold.
Then the Weyl function is the unique function, which satisfies the inequality
\begin{align}&      \label{2.20}
\int_0^{\infty}
\begin{bmatrix}
I_{m_1} & \vp(z)^*
\end{bmatrix}
u(x,z)^*u(x,z)
\begin{bmatrix}
I_{m_1} \\ \vp(z)
\end{bmatrix}dx< \infty .
\end{align}
\end{Cy}
\begin{proof}.  According to the equalities in \eqref{2.4} and \eqref{2.5}
and to the inequality \eqref{2.11} we derive
\begin{align}&      \label{2.21}
\int_0^{r}
\begin{bmatrix}
I_{m_1} & \vp(z)^*
\end{bmatrix}
u(x,z)^*u(x,z)
\begin{bmatrix}
I_{m_1} \\ \vp(z)
\end{bmatrix}dx =\frac{\I}{z - \ov z}\begin{bmatrix}
I_{m_1} & \vp(z)^*
\end{bmatrix}
\\ \nn &
\times \big(
\mf{A}(0,z)-\mf{A}(r,z)\big)
\begin{bmatrix}
I_{m_1} \\ \vp(z)
\end{bmatrix}\leq \frac{\I}{z - \ov z}\begin{bmatrix}
I_{m_1} & \vp(z)^*
\end{bmatrix}
\mf{A}(0,z)\begin{bmatrix}
I_{m_1} \\ \vp(z)
\end{bmatrix}.
\end{align}
Inequality \eqref{2.20} is immediate from \eqref{2.21}. Moreover,
as $u^*u\geq -\mf{A}$, the inequality \eqref{2.12} yields
\begin{align}&      \label{2.22}
\int_0^{r}
\begin{bmatrix}0 &
I_{m_2} 
\end{bmatrix}
u(x,z)^*u(x,z)
\begin{bmatrix}
0 \\ I_{m_2} 
\end{bmatrix}dx \geq 
r I_{m_2} .
\end{align}
By \eqref{2.22}, the function satisfying \eqref{2.20} is unique.
\end{proof}
\begin{Rk}\label{RkDefW} In view of Corollary \ref{CyW2}, inequality
\eqref{2.20} can be used as an equivalent definition of the Weyl function.
Definition of the form \eqref{2.20} is a more classical one and deals with
solutions of \eqref{1.1} which belong to $L^2(0, \, \infty)$. Compare with definitions
of Weyl-Titchmarsh or $M$-functions for discrete and continuous
systems in \cite{CG2, LS, Mar, SaA1, SaA2, SaL3, T1, T2} (see also references therein).
\end{Rk}
Our last proposition in this section is dedicated to a property of the Weyl
function,  the analog of which may be used as a definition of  generalized
Weyl functions in more complicated non-self-adjoint cases (see, e.g.,
\cite{FKS1, SaA2, SaA21}). 
\begin{Pn}\label{PnGW} Let Dirac system \eqref{1.1} on $[0, \, \infty)$
 be given, and assume that $v$ is  locally summable.
 Then,  the following inequality
\begin{align}&      \label{6.37}
\sup_{x \leq l, \, z\in \BC_+}\left\|\E^{-\I xz}u(x,z)\begin{bmatrix}
I_{m_1} \\ \vp(z)
\end{bmatrix}\right\|<\infty
\end{align}
holds on  any finite interval $[0, \, l]$
for the Weyl function $\vp$ of this system. 
\end{Pn}
\begin{proof}. We fix some $l$. Now, choose $x$ such that $0<x\leq l<\infty$. Because of \eqref{2.3} the Weyl function $\vp$ admits representations \eqref{2.2} 
(i.e., $\vp(z)=\vp(x,z, \clp)$). Hence, we can use
\eqref{2.1} and \eqref{2.9} to get
\begin{align}&      \label{6.38}
\Psi(x,z)^* j \Psi(x,z) \geq 0, \quad
\Psi(x,z):=\E^{-\I xz}u(x,z)\begin{bmatrix}
I_{m_1} \\ \vp(z)
\end{bmatrix}.
\end{align}
On the other hand, equation \eqref{1.1} and definition of $\Psi$
in \eqref{6.38} imply that
\begin{align}&      \nn
\frac{d}{dx}
\Big(\E^{-2xM}\Psi(x,z)^*(I_m+ j) \Psi(x,z) \Big)
=\E^{-2xM}\Psi(x,z)^*
\\ &      \label{6.39}
\times  \Big(\I \big((I_m+ j)jV-Vj(I_m+ j)\big)-2M(I_m+ j)\Big) \Psi(x,z)
\\ & \nn =2\E^{-2xM}\Psi(x,z)^*\begin{bmatrix}
-2 M I_{m_1} & \I v(x)\\ - \I v(x)^* &0
\end{bmatrix}\Psi(x,z), \quad
M:=\sup_{x<l}\|V(x)\|.
\end{align}
Using \eqref{6.38} and \eqref{6.39} we derive
\begin{align}&      \nn
\frac{d}{dx}
\Big(\E^{-2xM}\Psi(x,z)^*(I_m+ j) \Psi(x,z) \Big) \leq
2\E^{-2xM}\Psi(x,z)^*
\\ & \label{6.40}
\times \left(\begin{bmatrix}
 & \I v(x)\\ - \I v(x)^* &0
\end{bmatrix}-MI_m \right)\Psi(x,z)\leq 0.
\end{align}
Finally, inequalities  \eqref{6.38} and \eqref{6.40}  lead us to
\begin{align}&      \label{6.41}
\Psi(x,z)^* \Psi(x,z) \leq
\Psi(x,z)^*(I_m+ j) \Psi(x,z)  \leq
2\E^{2xM}I_{m_1},
\end{align}
and \eqref{6.37} follows.
\end{proof}
\section{Direct and inverse problem: explicit solutions} \label{dies}
\setcounter{equation}{0}
Various versions of B\"acklund-Darboux transformations are actively
used to construct explicit solutions of linear and integrable nonlinear
equations (see, e.g., \cite{Ci, GeH, Gu, Mar2, MS, ALS10, ZM} and numerous
references therein). For the spectral and scattering results that follow
from B\"acklund-Darboux transformations and related commutation and factorization methods see, for instance, publications \cite{cr, de, FKS0, fg, gt, GKS6, KoSaTe,
kr57, SaA7}. Here we will give explicit solutions of our  direct and inverse problems using the GBDT version of the B\"acklund-Darboux transformation
(see \cite{FKRS, FKS0, GKS6, ALS94, SaA7, ALS10} and references therein).

To obtain explicit solutions, we consider
$m_1 \times m_2$
potentials $v$ of the form
\begin{align} \label{7.1}&
v(x)=-2i \vt _{1}^{\, *}\E^{ix \alpha^{*}} \S
(x)^{-1}\E^{ix
\alpha }\vt_2,
\end{align}
where some  $n\in \BN$ is fixed and
the $n \times n$ matrix function $\S$ 
is given by
the formula
\begin{align}   \label{7.2}&
\S(x)=\S_0+ \int_{0}^{x} \Lambda(t) \Lambda (t)^{*}dt
\quad(\S_0>0),
\quad
\Lambda (x)= \begin{bmatrix}  \E^{-ix \alpha } \vt_{1} 
& \E^{ix
\alpha } \vt_2 \end{bmatrix}.
\end{align}
Here $\a$, $\vt_1$, and $\vt_2$ are $n \times n$, $n \times
m_1$, and $n
\times m_2$ parameter matrices,  and the following
matrix identity holds:
\begin{align} \label{7.3}&
\a \S_0-\S_0 \a^*=\I (\vt_1 \vt_1^*- \vt_2 \vt_2^*).
\end{align}
Clearly, $\S(x)$ is invertible for $x\geq 0$ and the potential $v$ in \eqref{7.1}
is well-defined. 
\begin{Dn}\label{Dnpsex}
The $m_1 \times m_2$
potentials $v$ of the form \eqref{7.1}, where relations \eqref{7.2} and \eqref{7.3}
hold, are called generalized pseudo-exponential potentials. It is said that
$v$ is generated by the parameter matrices $\a$, $\S_0$, $\vt_1$, and $\vt_2$.
\end{Dn}
According to \cite[Theorem 3]{ALS94}
(see also \cite{FKS0}), the fundamental solution $u$ of system \eqref{1.1}, where
$V$ is given by \eqref{1.2}, $v$   is a generalized pseudo-exponential potential, and $u$ is
normalized by \eqref{1.3}, admits representation
\begin{align} \label{7.4}&
u(x, z)=w_{\a}(x, z)\E^{ix z j}w_{\a}(0, z)^{-1}.
\end{align}
Here we have
\begin{align} \label{7.5}&
w_{\a}(x, z):=I_m+ij \Lam(x)^*\S(x)^{-1}(z I_n- \a)^{-1}
\Lam(x).
\end{align}

Note that the case  $m_1=m_2$ (i.e., the case of the 
pseudo-exponential potentials) was treated in greater detail in \cite{GKS6}
(see \cite{GKS6} and references therein for the term {\it pseudo-exponential},
itself, too).  Formulas (\ref{7.2}) and (\ref{7.3}) yield
\begin{equation} \label{7.6}
\a \S(x)-\S(x) \a^*=i \Lam (x)j \Lam (x)^*.
\end{equation}
Identity \eqref{7.6}, in its turn, implies that $w_{\a}(z)$  
is a transfer matrix function
in Lev Sakhnovich form \cite{SaL1, SaL2, SaL2', SaL3}.
However, $w_{\a}(x, z)$  possesses an additional variable
$x$ and the way, in which this matrix function depends on $x$,
is essential. 

From \cite[formula (2.9)]{FKS0}, where $W_{11}$ and $W_{21}$
are $m_1 \times m_1$ and $m_2 \times m_1$ blocks of $w_{\a}(0, z)$:
\begin{align} \label{7.7}&
w_{\a}(0, z)=:\{W_{ij}(z)\}_{i,j=1}^2,
\end{align}
we see that
\begin{align} \label{7.8}&
W_{21}(z)W_{11}(z)^{-1}= - i\vt_2^*\S_0^{-1}(z I_n
- \t)^{-1}
\vt_1 , \quad \t:=\a-i\vt_1\vt_1^*\S_0^{-1}.
\end{align}
We note that  \cite[formulas (2.6)  and (2.7)]{FKS0} imply that $W_{11}(z)$ 
is always well-defined and invertible for $z\not\in \s(\a)\cup\s(\t)$, where
$\s$ denotes spectrum.

Relations \eqref{7.4}, \eqref{7.7}, and \eqref{7.8} are basic to solve the direct
problem for Dirac systems with the generalized pseudo-exponential potentials
\eqref{7.1}.
\begin{Tm}\label{TmDPes} Let Dirac system \eqref{1.1} on $[0, \, \infty)$
 be given and assume that $v$ is  a generalized pseudo-exponential potential,
 which is generated by the matrices $\a$, $\S_0$, $\vt_1$, and $\vt_2$.
 Then the Weyl function $\vp$ of  system \eqref{1.1} has the form:
\begin{align} \label{7.9}&
\vp(z)=- i\vt_2^*\S_0^{-1}(z I_n
- \t)^{-1}
\vt_1 , \quad \t=\a-i\vt_1\vt_1^*\S_0^{-1}.
\end{align}
\end{Tm}
\begin{proof}. We compare \eqref{7.8} and \eqref{7.9} to see that
\begin{align} \label{7.10}&
\vp(z)=W_{21}(z)W_{11}(z)^{-1}.
\end{align}
Because of \eqref{7.4}, \eqref{7.7}, and \eqref{7.10} we have
\begin{align} \label{7.11}&
u(x,z)
\begin{bmatrix}
I_{m_1} \\ \vp(z)
\end{bmatrix}=\E^{ixz}w_{\a}(x, z)\begin{bmatrix}
I_{m_1} \\ 0
\end{bmatrix}W_{11}(z)^{-1}.
\end{align}
To consider the matrix function $\Lam^*\S^{-1}$, which appears in the definition \eqref{7.5} of $w_{\a}$,
we derive from \eqref{7.2} that
\begin{align} \label{7.12}&
\S(x)^{-1}\Lam(x)\Lam(x)^*\S(x)^{-1}= -\frac{d}{dx}\S(x)^{-1}.
\end{align}
It is immediate also from \eqref{7.2} that $\S(x)>0$. 
Therefore, using  \eqref{7.12} we get 
\begin{align} \label{7.13}&
\int_0^{\infty} \S(t)^{-1}\Lam(t)\Lam(t)^*\S(t)^{-1}dt\leq \S_0^{-1}.
\end{align}
Furthermore, the last equality in \eqref{7.2} implies that
\begin{align} \label{7.14}&
\sup_{\Im z>\|\a\|+\ve }\|\E^{ixz}\Lam(x)\|<M_{\ve} \quad (\ve>0).
\end{align}
It follows from \eqref{7.5}, \eqref{7.13}, and \eqref{7.14} that
the entries of the right-hand side of \eqref{7.11} are well-defined and
uniformly bounded in the $L^2(0,\, \infty)$ norm with respect to $x$
for all $z$ such that $\Im z \geq \max \big(\|\a\|, \,  \|\t\|\big)+\ve$ and $\ve>0$.
Hence, taking into account \eqref{7.11} we see that \eqref{2.20} holds for $z$
from the  mentioned above domain. So, according to the
uniqueness statement in
Corollary \ref{CyW2},
$\vp(z)$ of the form \eqref{7.9} coincides with the Weyl function in that domain. Since the Weyl function
is analytic in $\BC_+$, the matrix function $\vp$ coincides with it in $\BC_+$
(i.e., $\vp$ is the Weyl function, indeed).
\end{proof}
For the case that $v$ is a  
generalized pseudo-exponential potential, where $\S_0>0$, our Weyl function 
coincides with the reflection coefficient from \cite{FKS0} (see 
\cite[Theorem 3.3]{FKS0}). Hence, the solution of our inverse problem
can be considered as a particular case of  the solution of the inverse problem
from \cite[Theorem 4.1]{FKS0}, where the singular case $\S_0\not>0$ was
treated too.

Before we formulate the procedure to solve inverse problem, some 
results  on rational matrix functions and notions from system and  control theories are required (see, e.g., \cite{KFA, LR}). Let $\vp(z)$ be a strictly proper rational matrix function, that is, such a rational matrix function that
\begin{align} \label{7.15}&
\lim_{z\to \infty} \vp(z)=0.
\end{align}
Then $\vp$ admits representations (also called realizations):
\begin{align} \label{7.16}&
 \vp(z)=C_N(z I_N-\cla_N)^{-1}B_N,
\end{align}
where $C_N$, $\cla_N$, and $B_N$ are $m_2 \times N$,
$N \times N$, and $N \times m_1 $, respectively, matrices. Here $N\in \BN$,
and $m_1$ ($m_2$) denotes the number of columns (rows) of $\vp$.
\begin{Dn}\label{DnST}
The minimal possible value of $N$ in realizations \eqref{7.16} is called
the McMillan degree of $\vp$, and we denote this value by $n$.
Realizations \eqref{7.16}, where $N=n$, are called the minimal realizations.
\end{Dn}
From \cite[Theorems 21.1.3, 21.2.1]{LR} we easily see that for  a minimal
realization
\begin{align} \label{7.17}&
 \vp(z)=C(z I_n-\cla)^{-1}B
\end{align}
of a matrix $\vp$, which is non-expansive on $\BR$ and has no poles
in $\BC_+$, there is a positive solution $X>0$ of the Riccati equation 
\begin{align} \label{7.18}&
XC^*CX+\I(X\cla^*-\cla X)+BB^*=0.
\end{align}
Furthermore, all the hermitian
solutions of \eqref{7.18} are positive.
\begin{Tm}\label{TmIpes} Let $\vp(z)$ 
be a strictly proper rational matrix function, which  is non-expansive on $\BR$ and has no poles in $\BC_+$.  Assume that \eqref{7.17} is its minimal realization
and that $X>0$ is a solution of \eqref{7.18}. 

Then $\vp(z)$ is the Weyl function of Dirac system, the potential of which
is given by \eqref{7.1} and \eqref{7.2}, where
\begin{align} \label{7.19}&
\a=\cla +\I BB^*X^{-1}, \quad \S_0=X, \quad \vt_1=B, \quad \vt_2=-\I X C^*.
\end{align}
This solution of the inverse problem is unique in the class of Dirac systems
with locally bounded potentials.
\end{Tm}
\begin{proof}.  From \eqref{7.19} we see that
\[
\a\S_0-\S_0\a^*=\cla X-X\cla^*+2iBB^*, \quad
\I(\vt_1\vt_1^*-\vt_2\vt_2^*)=\I BB^*-\I XC^*CX,
\]
and so \eqref{7.3} is equivalent to \eqref{7.18}. Since \eqref{7.3} holds,
we apply Theorem \ref{TmDPes}.   Theorem \ref{TmDPes} implies that the Weyl function of the Dirac system, where $v$ is given by \eqref{7.1},
has the form \eqref{7.9}.
Next, we substitute \eqref{7.19} into \eqref{7.9}, to derive that the right-hand
sides of \eqref{7.17} and the first equality in \eqref{7.9} coincide.
In other words, the Weyl function of our system admits representation
\eqref{7.17}.

Finally, the uniqueness of the solution of the inverse problem
follows from Theorem \cite[Theorem 4.1]{FKRS3}.
\end{proof}
We note that the corresponding uniqueness result in \cite{FKS0}
was proved only for the class of systems with the generalized
pseudo-exponential potentials.

Because of the second equality in  \eqref{7.8} and identity
\eqref{7.3}, the matrix $\t$ satisfies another identity:
$\t \S_0-\S_0 \t^*=-\I (\vt_1 \vt_1^*+ \vt_2 \vt_2^*)$, that is,
\begin{align} \label{7.22}&
 \S_0^{-1}\t -\t^* \S_0^{-1}=-\I \S_0^{-1} (\vt_1 \vt_1^*+ \vt_2 \vt_2^*)\S_0^{-1}.
\end{align}
If $f\not=0$ is an eigenvector of $\t$ (i.e., $\t f=\la f$), identity \eqref{7.22}
implies that
\begin{align}  \label{7.23}&
(\la - \ov{\la})f^*\S_0^{-1}f=-\I f^*\S_0^{-1} (\vt_1 \vt_1^*+ \vt_2 \vt_2^*)\S_0^{-1}f.
\end{align}
Since $\S_0>0$, we derive from \eqref{7.23} that
\begin{align}  \label{7.24}&
\s(\t) \subset \BC_- \cup \BR .
\end{align}

Real eigenvalues of $\t$ play a special role in the spectral
theory of  an operator, which corresponds to the
Dirac system with a generalized pseudo-exponential potential
(see, e.g., \cite{GKS5} for the case of square potentials).
In our case the operator $\clh$  corresponding to the
Dirac system is defined in a way, which is similar
to  the definition from \cite{GKS5}, but the initial condition is
quite different. Namely, we determine $\clh$ 
by the differential expression
\begin{align} \label{7.20}&
\clh_{de} y=-\I j \frac{d}{dx}y-Vy,
\end{align}
and by its domain $\cld(\clh)$, which consists of all locally absolutely
continuous $\BC^m$-valued functions $y$ in $L^2_m(0, \, \infty)$, 
such that 
\begin{align} \label{7.21}&
\clh_{de} y\in L^2_m(0, \, \infty), \qquad y(0)=0.
\end{align}
\begin{Pn}\label{PnEV} Let the conditions of Theorem \ref{TmDPes}
hold,  let $\t$ be given by the second relation in \eqref{7.8}, and let
$\la$ be an eigenvalue of $\t$:
\begin{align} \label{7.25}&
\t f=\la f, \quad f \not= 0, \quad \la \in \BR .
\end{align}
Then, the matrix function
\begin{align} \label{7.26}&
g(x):=j\Lam(x)^*\S(x)^{-1}f
\end{align}
is a bounded state of $\clh$ and $\clh g=\la g$.
\end{Pn}
\begin{proof}.  First, we show that formulas \eqref{7.23} and \eqref{7.25} yield
\begin{align} \label{7.27}&
 \vt_1^*\S_0^{-1}f=0, \quad    \vt_2^*\S_0^{-1}f=0, \quad          \a f=\la f.
 \end{align}
Indeed, the first two equalities in \eqref{7.27} easily follow from
\eqref{7.23} for the case that $\la =\ov \la$. The equality $ \a f=\la f$
is immediate from $\t f = \la f$, definition of $\t$ in \eqref{7.8}, and equality
$\vt_1^*\S_0^{-1}f=0$.

Next, we show that
\begin{align} \label{7.28}&
\big(j\Lam^*\S^{-1}\big)^{\prime}=\I j^2 \Lam^*\S^{-1}\a+
\big(\Lam^*\S^{-1}\Lam - j \Lam^*\S^{-1}\Lam j\big)j \Lam^*\S^{-1}.
\end{align}
Formula \eqref{7.28} follows from  a general GBDT formula
\cite[(3.14)]{ALS10} and also from its Dirac system subcase 
\cite[(2.13)]{ALS10}, but it will be convenient to prove \eqref{7.28} directly.
We note that  formula \eqref{7.2} implies
\begin{align} \label{7.29}&
\Lam^{\prime}= -\I \a \Lam j, \qquad \S^{\prime}=\Lam\Lam^*,
\end{align}
and formula \eqref{7.6} can be rewritten as
\begin{align} \label{7.30}&
\a^* \S^{-1}=\S^{-1}\a- \I \S^{-1}\Lam j \Lam^*\S^{-1}.
\end{align}
Since $j^2=I_m$, using \eqref{7.29} and \eqref{7.30} we obtain \eqref{7.28}.

Now, partitioning $\Lam$ into two blocks and using \eqref{7.1} and \eqref{7.2}, we see that
\begin{align} \label{7.31}&
v(x)=-2\I \Lam_1(x)^*\S(x)^{-1}\Lam_2(x), \quad \Lam=: \begin{bmatrix} \Lam_1 & \Lam_2 \end{bmatrix}.
\end{align}
In view of \eqref{1.2} and  \eqref{7.31} we have
\begin{align} \label{7.32}&
\Lam^*\S^{-1}\Lam - j \Lam^*\S^{-1}\Lam j=\I j V.
\end{align}
Applying both sides of \eqref{7.28} to $f$ and taking into account the last equality in \eqref{7.27} and relation \eqref{7.32},
we derive
\begin{align} \label{7.33}&
\big(j\Lam(x)^*\S(x)^{-1}f\big)^{\prime}=\I \la j^2 \Lam(x)^*\S(x)^{-1}f+
\I jV(x)j \Lam(x)^*\S(x)^{-1}f.
\end{align}
Because of  \eqref{7.20} and  \eqref{7.26}, we can rewrite \eqref{7.33} as
\begin{align} \label{7.34}&
\clh_{de}g=\la g,
\end{align}
and it remains to show that $g\in \cld(\clh)$, that is, that $g \in L^2_m(0,\, \infty)$ and \eqref{7.21}
holds for $y=g$. From  \eqref{7.13} and  \eqref{7.34} we see that $g, \, \clh_{de}g \in L^2_m(0,\, \infty)$.
Finally, the initial condition 
\begin{align} \label{7.35}&
g(0)=\begin{bmatrix} \vt_1 & - \vt_2 \end{bmatrix}^*\S_0^{-1}f=0
\end{align}
 is immediate from \eqref{7.2}, \eqref{7.26}, and \eqref{7.27}.
\end{proof}


{\bf Acknowledgement.}
The work of I.Ya. Roitberg was supported by the 
German Research Foundation (DFG) under grant no. KI 760/3-1 and
the work of A.L. Sakhnovich was supported by the Austrian Science Fund (FWF) under Grant  no. Y330.



\end{document}